1. 2(x+1)=3x-1.
2. 6(x+1)=12(x-3).

Answer by elima(1433)

1. 2(x+1)=3x-1.
First distribute the 2;
2x+2=3x-1
subtract 3x and 2 from each side;
2x-3x=-1-2
-x=-3
x=3
2. 6(x+1)=12(x-3).
Distribute the 6 and 12;
6x+6=12x-36
subtract 12x and add 6 to each side;
6x-12x=-36+6
-6x=-30
x=5
Hope this helps
=)